In a recent survey of 205 never-married single women, 78% thought they would eventually find, and marry, their perfect mate. The standard error for a confidence interval for this proportion is

.4142

.0008

.1716

.0289
Answer #1

Solution :

Given that,

n = 205

Point estimate = sample proportion = \hat p = 0.78

1 - \hat p = 1 - 0.78=0.22

SE= \sqrt ((\hat p * (1 - \hat p )) / n)

=(\sqrt((0.78*0.22) / 205)

SE =0.0289